Greta, an elderly investor, has a degree of risk aversion of A = 3 when applied to return on wealth over a one-year horizon. She is pondering two portfolios, the S&P 500 and a hedge fund, as well as a number of one-year strategies. (All rates are annual and continuously compounded.) The S&P 500 risk premium is estimated at 6.6% per year, with a SD of 21.6%. The hedge fund risk premium is estimated at 11.6% with a SD of 36.6%. The returns on both of these portfolios in any particular year are uncorrelated with its own returns in other years. They are also uncorrelated with the returns of the other portfolio in other years. The hedge fund claims the correlation coefficient between the annual returns on the S&P 500 and the hedge fund in the same year is zero, but Greta is not fully convinced by this claim. Calculate Greta’s capital allocation using an annual correlation of 0.3.(Do not round your intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

Risk free rate 1 year Treasury rate 2.00% (Assumed) Fund S&P 500 (S) Hedge Fund (H) Risk Premium Return 6.60% 11.60% Standard

49 Portfolio variance is given by following formula Portfolio variance *(R o (R)2(wA) (Wa)*Cov(A, B) Where wA and wB are weigFormula sheet

A B C D E F G H I J
2
3
4
5 Risk free rate = 1 year Treasury rate 0.02 (Assumed)
6
7 Fund Risk Premium Return Standard Return
8 S&P 500 (S) 0.066 =D8+D5 0.216
9 Hedge Fund (H) 0.116 =D9+D5 0.366
10
11 Correlation 0.3
12
13 Cov (S,H) =?*?S*?B
14 =D11*F8*F9 =D11*E8*E9
15
16 Calculation of weight of optimal portfolio:
17 The minimum variance portfolio of two Risky asset can be calculated as follows:
18
19
20
21
22
23
24
25 Optimal asset allocation in S&P 500, wmin(S) =(F9^2-D14)/(F9^2+F8^2-2*D14)
26 Optimal asset allocation in Hedge Fund, wmin(H) =1-D25
27
28 Calculation of expected return of minimum variance portfolio:
29 Portfolio expected return can be calculated as follows:
30
31
32
33
34 Where wi and ri are the weights and return of assets Ai
35
36 Expected return of minimum variance portfolio =Sum Product of portfolio weight and return
37 =SUMPRODUCT(D25:D26,E8:E9) =SUMPRODUCT(D25:D26,E8:E9)
38
39 Expected return of minimum variance portfolio =D37
40
41
42 Given the risk aversion (A) of the investor, the weight of risky assets in the portfolio is given by following equation:  
43
44
45
46 Where E(rp) is the expected return of the risky portfolio, rf is the risk free rate,
47 A is risk aversion of investor and ?p2 is the variance of risky portfolio.
48
49
50 Portfolio variance is given by following formula:
51 Portfolio variance = w2A*?2(RA) + w2B*?2(RB) + 2*(wA)*(wB)*Cov(A, B)
52 Where wA and wB are weights of assets A and B, ?A and ?B are standard deviation of assets A and B.
53 Given the following data:
54 S&P 500 (S) Hedge Fund (H)
55 Expected Return =E8 =E9
56 Standard Deviation =F8 =F9
57 Weight =D25 =D26
58 Covariance =D14
59
60 Variance of Risky portfolio =w2A*?2(RA) + w2B*?2(RB) + 2*(wA)*(wB)*Cov(A, B)
61 =(D57*D56)^2+(E57*E56)^2+2*D57*E57*D58 =(D57*D56)^2+(E57*E56)^2+2*D57*E57*D58
62
63 Hence Variance of risky portfolio is =D61
64
65 Calculation of weight of risky asset in the portfolio:
66 Given the risk aversion (A) of the investor, the weight of risky assets in the portfolio is given by following equation:  
67
68
69
70 Where E(rp) is the expected return of the risky portfolio, rf is the risk free rate,
71 A is risk aversion of investor and ?p2 is the variance of risky portfolio.
72 Using the following data:
73 Expected return of risky portfolio =D39
74 risk free rate =D5
75 Risk Aversion 3
76 Variance of risky portfolio =D63
77
78 Weight of risky assets in the portfolio can be calculated as follows:
79 Weight of risky assets in the portfolio =(E(rp)-rf)/(A*?p2)
80 =(D73-D74)/(D75*D76) =(D73-D74)/(D75*D76)
81
82 Hence Weight of risky assets in the portfolio =D80
83 Weight of risk free asset in the portfolio =1-Weight of risky asset in the portfolio
84 =1-D82
85
86 Using the weight of risky asset in the portfolio and the weight of individual fund in the risky asset,
87 weight of fund in the portfolio can be calculated as follows:
88 Using the following data:
89 Weight of risky assets in the portfolio =D82
90 Weight of S&P 500 in risky asset =D25
91 Weight of Hedge fund in risky asset =D26
92
93 Weight of S&P 500 in the portfolio =Weight of risky asset in the portfolio*weight of S&P 500 in risky asset
94 =D89*D90 =D89*D90
95
96 Weight of Hedge fund in the portfolio =Weight of risky asset in the portfolio*weight of Hedge fund in risky asset
97 =D89*D91 =D89*D91
98
99 Hence Capital allocation is as follows:
100 Weight of risk free asset =D84
101 Weight if S&P 500 =D94
102 Weight of Hedge fund =D97
103